QUESTION9 rice noor in this market at $3.00, what will be the total loss in welfare to the QD.64-16P and s-16ρ-Rif the government imposes a Assume that the market for Good X defined as follows economy? 24.50 549.00 $9.00 $8.00 There is no loss because producers receive a higher price for their product
Answer #1

Answer: 3rd option; $9.00

Market equilibrium price and quantity should be calculated first.

Condition: QD = QS

64 – 16P = 16P – 8

32P = 64 + 8

32P = 72

P = 72/32 = 2.25

By putting this value in either QD or QS

QD = 64 – 16P

       = 64 – 16 × 2.25

       = 64 – 36

       = 28 = QS

Equilibrium price is $2.25 and quantity (Q) is 28 units.

Now at $3 price QD would be as below,

QD = 64 – 16P

       = 64 – 16 × 3

       = 64 – 48

       = 16

Loss of QD = 0.5 × Difference in QD × Difference in price

                    = 0.5 × (28 – 16) × (3 – 2.25)

                    = 0.5 × 12 × 0.75

                    = 4.5

Now at 16 units, supply price would be as below,

QS = 16P – 8

16 = 16P – 8

16P = 16 + 8

P = 24/16 = 1.5

Loss of QS = 0.5 × Difference in QD × Difference in supply price

                    = 0.5 × (28 – 16) × (2.25 – 1.5)

                    = 0.5 × 12 × 0.75

                    = 4.5

Total loss = Loss of QD + Loss of QS

                 = 4.5 + 4.5

                   = $9